### v1

# Cancellation of phonon hopping in trapped ions by modulation of the trap potential

Abstract
The local modes of trapped ions can be used to construct an analog quantum simulator and a digital quantum computer. However, the control of the phonon hopping remains difficult because it proceeds among all the local modes through the Coulomb coupling. We propose a method to cancel the phonon hopping among a given set of local modes by applying a sequence of phase shift gates implemented through the modulation of the trap potential. We analyze the error scaling in the algorithm to treat three or more modes and show that the error can be suppressed by repeating the pulse sequence. The duration of the phase shift gate in the present method can be as short as a few microseconds, which is an order of magnitude faster than the laser-based method. This short duration of the phase shift gate facilitates the suppression of the gate error. The present method can be applied to the simulation of bosonic systems as well as to the continuous variable encoding of quantum computing using trapped ions.
pacs:
preprint:
I INTRODUCTION
The motional oscillator of trapped ions in a harmonic potential provides a controllable system of phonons, which can be used for continuous variable (CV) encoding of quantum information processing [1, 2] and quantum simulation of bosonic systems [3, 4, 5]. In general, the Hamiltonian of the system composed of harmonic oscillators can be described either by the local mode or by the collective mode. For the local mode, the Hamiltonian is composed of the on-site phonon interaction and the phonon hopping induced by the Coulomb repulsion, while for the collective mode, the Hamiltonian matrix is obtained by diagonalizing the local-mode Hamiltonian matrix.
The local mode scheme can be used to realize the quantum simulation of, e.g., a Bose–Hubbard model [3] and a Jaynes–Cummings–Hubbard model [4], and can also be used to implement a quantum walk [6] and a boson sampling [7]. Moreover, it has been shown that a logical qubit in the CV encoding called the Gottesman–Kitaev–Preskill (GKP) state can be constructed using a local mode [8, 9, 10].
However, in the local mode scheme, the manipulation of entanglement among two or more local modes remains difficult because the phonon hopping, a source of entanglement, originating from the Coulomb interaction cannot be turned off. On the other hand, in the collective mode scheme [11], the modes can be entangled using a beam-splitter gate, which can be switched on only by irradiating the ions with laser [11] or by applying the spatially and temporally varying electric potential [12].
There are several methods for the control of the phonon hopping. In the explanation below, we refer to a system having two local modes for simplicity, but the discussion can be extended to a system having more than two local modes. The first method is called the phonon blockade [13], where two local modes become off-resonant by the irradiation of one of the two local modes with laser light so that the phonon hopping can be suppressed. The method protects the phonon state of one mode, but disturb that of the other mode being excited by the laser light.
In the second method [14], the two ions are stored in distant traps where the phonon hopping is negligible, and they are transported into proximity by the deformation of the trap potential, kept for a certain time so that a desired amount of phonon hopping is induced, and then separated apart again. This scheme requires a large-size ion storage area in which the distance between the two ions can be sufficiently long.
The third method is based on the idea of the dynamical decoupling (DD) [7, 15]. In this method, two modes are resonant and the phonon hopping proceeds from $t=0$ to $T/2$, then an instantaneous $\pi$-phase shift $P$ is applied to one mode, which flips the sign of the phonon hopping Hamiltonian $H_{C}$, i.e., $H_{C}P=-PH_{C}$. Therefore, the time evolution from $t=T/2$ to $T$ cancels with the phonon hopping in the first half period. Because the phonon hopping proceeds even during the application of $P$, it is crucial to make the duration of $P$ negligibly short compared to the time scale of the phonon hopping.
In the original idea of the DD proposed by Shen et al.[7], the off-resonant sideband pulse was adopted to implement $P$. However, when the practical experimental conditions are considered [15], it is difficult to make the duration of $P$ sufficiently short because the rate of the off-resonant interaction is comparable to the hopping rate . Ohira et al.[15] showed that a shorter $\pi$-phase shift $P$ can be realized by adopting the resonant sideband pulse, but their scheme can be used only when the phonon number is 0 or 1.
In the present study, we adopt the third method and propose to implement the $\pi$-phase shift by modulating the trap potential using an electric pulse applied to the trap electrodes and show that the fidelity of the DD can be improved. The duration of $P$ can be a few microseconds, which is shorter by an order of magnitude than that can be achieved by the resonant sideband pulse approach. It should be noted that the present method can be applied to a system having any number of phonons. We also show that the original DD approach  [7] needs to be modified when the number of modes becomes larger than two even when the duration of $P$ is negligibly small.
II Method
II.1 A sequence of phase shifts for the dynamical decoupling
We consider a chain of ions and assume that the frequencies of the oscillation along the radial direction $\tilde{x}$, perpendicular to the trap axis, are a few MHz while the hopping rate is a few kHz. This condition allows us to adopt the local-mode picture along $\tilde{x}$ axis. Therefore, if the number of trapped ions is $M$, we consider $M$ local modes and denote the set of those modes as $\mathcal{U}=\{q_{j};0\leq j\leq M-1\}$. The scaled coordinate of the $j$-th mode $x_{j}$ is represented using the annihilation and creation operators as $x_{j}=(a_{j}^{\dagger}+a_{j})/\sqrt{2}$. As shown in Appendix. A, when the distance between adjacent ions is much larger than the displacement of the respective ions from their equilibrium positions, the Hamiltonian is composed of the on-site interaction $H_{0}$ and the Coulomb interaction $H_{C}$ as
$H$ $=H_{0}+H_{C}$ (1)
$H_{0}$ $=\sum_{j=0}^{M-1}\hbar\omega_{0}\left(a_{j}^{\dagger}a_{j}+\frac{1}{2}\right)$ (2)
$H_{C}$ $=\sum_{j>k}\frac{\hbar\kappa_{j,k}}{2}\left(a_{j}^{\dagger}+a_{k}^{\dagger}\right)\left(a_{j}+a_{k}\right)$
$\simeq\sum_{j>k}\frac{\hbar\kappa_{j,k}}{2}\left(a_{j}^{\dagger}a_{k}+a_{j}a_{k}^{\dagger}\right)=\sum_{j>k}H_{j,k},$ (3)
where $\omega_{0}$ in Eq. (2) is the secular frequency obtained under the influance of the Coulomb interactions. The rotating-wave approximation (RWA) has been applied to $H_{C}$ so that the fast oscillating terms such as $a_{j}a_{k}+a_{j}^{\dagger}a_{k}^{\dagger}$ can be omitted. The Hamiltonian in the interaction picture with respect to $H_{0}$ is obtained as $H_{I}=H_{C}$. The time evolution under $H_{I}$ for the time duration $\tau$ is given by
$\exp(-\mathrm{i}\tau H_{I}/\hbar)$
$=B_{M}(\bm{\theta}_{\tau})\equiv\exp\left(-\mathrm{i}\tau\sum_{j>k}\frac{\kappa_{j,k}}{2}\left(a_{j}^{\dagger}a_{k}+a_{j}a_{k}^{\dagger}\right)\right),$ (4)
where $B_{M}$ is an $M$-mode entangling gate and $(\bm{\theta}_{\tau})_{j,k}\equiv\tau\kappa_{j,k}/2$.
Because $B_{M}$ induces the phonon hopping among all the modes, when we need to induce the phonon hopping among a set of modes $\mathcal{S}$ as
$B_{\mathcal{S}}=\exp\left(-\mathrm{i}\tau\sum_{\begin{subarray}{c}j>k\\ j,k\in\mathcal{S}\end{subarray}}\frac{\kappa_{j,k}}{2}\left(a_{j}^{\dagger}a_{k}+a_{j}a_{k}^{\dagger}\right)\right),$ (5)
we should cancel the phonon hopping among the other modes $q_{r}\in\mathcal{S}^{c}$ as well as that between $q_{r}$ and $q_{s}$ for $q_{s}\in\mathcal{S}$ and $q_{r}\in\mathcal{S}^{c}$. For example, in order to implement the 50:50 beam-splitter between $q_{j}$ and $q_{j+1}$, we need to wait for $\tau=T_{\mathrm{50:50}}=(\pi/4)/(\kappa_{j+1,j}/2)$ while the phonon hopping involving $q_{r}\in\mathcal{S}^{c}$ should be cancelled, i.e., the phonon hopping between $q_{r_{1}}$ and $q_{r_{2}}$ for $q_{r_{1}},q_{r_{2}}\in\mathcal{S}^{c}$, that between $q_{j}$ and $q_{r}$, and that between $q_{j+1}$ and $q_{r}$ should be cancelled. Because the waiting time $\tau$ depends on the operation time of the gate applied to the modes in $\mathcal{S}$, we consider the DD to cancel the phonon hopping occurring in the course of the time propagation from $t=0$ to arbitrary time $T$. First, we explain the basic idea of the DD for a two mode case, and then extend the method to a general case with the larger number of modes.
II.1.1 Two-mode DD
When only two modes are involved, the time propagation for the time duration $\tau$ corresponds to a two-mode beam-splitter $B^{\mathrm{TM}}_{1,0}(\theta_{\tau})$ given by
$B_{2}(\theta_{\tau})$
$=B^{\mathrm{TM}}_{1,0}(\theta_{\tau})=\exp\left(-\mathrm{i}\tau\frac{\kappa_{1,0}}{2}\left(a_{1}^{\dagger}a_{0}+a_{1}a_{0}^{\dagger}\right)\right),$ (6)
where $\theta_{\tau}\equiv\tau\kappa_{1,0}/2$. As shown in Appendix B, when a $\pi$-phase shift is defined by
$P_{j}=\exp\left(-\mathrm{i}\pi a_{j}^{\dagger}a_{j}\right),~{}j=0~{}\mathrm{or}~{}1,$ (7)
the sign of $H_{I}$ flips as $H_{I}P_{j}=-P_{j}H_{I}$, from which $B^{\mathrm{TM}}_{1,0}(\theta)P_{j}=P_{j}B^{\mathrm{TM}}_{1,0}(-\theta)$ can be derived. Therefore, the phonon hopping occurring in the course of the time propagation from $t=0$ to $T$ represented by $B^{\mathrm{TM}}_{1,0}(\theta_{T})$ can be canceled by applying either $P_{0}$ or $P_{1}$ at $t=T/2$ and $T$ as
$U_{\mathrm{DD}}=P_{j}B^{\mathrm{TM}}_{1,0}(\theta_{T/2})P_{j}B^{\mathrm{TM}}_{1,0}(\theta_{T/2})=I,\,j=0~{}\mathrm{or}~{}1.$ (8)
where $I$ denotes the identity operator. It should be noted that we assume here the duration of $P_{j}$ is negligibly short. We can interpret Eq. (8) as the following sequence; (i) the phonon hopping proceeds for the first half period ($B^{\mathrm{TM}}_{1,0}(\theta_{T/2})$), (ii) $P_{j}$ is applied to flip the sign of $H_{I}$, (iii) the phonon hopping proceeds with $-H_{I}$ in the latter half, which cancels the phonon hopping in the first half as
$B^{\mathrm{TM}}_{1,0}(\theta_{T/2})P_{j}B^{\mathrm{TM}}_{1,0}(\theta_{T/2})$
$=P_{j}B^{\mathrm{TM}}_{1,0}(-\theta_{T/2})B^{\mathrm{TM}}_{1,0}(\theta_{T/2})=P_{j}$ (9)
and (iv) the $\pi$-phase shift is compensated by the second $P_{j}$ at the end by using $P_{j}^{2}=I$.
II.1.2 DD for three or more modes
The DD for the general case with more than three modes, called the concatenation, was proposed in Ref. [7]. Here, we give an algorithm to construct a pulse sequence for the concatenation. In the first step, we consider a chain of $M$ modes and group them into two sets. Although the DD works for an arbitrary grouping, we choose for the simplicity the two sets as the first half $S_{0}=\{q_{j};~{}0\leq j\leq\lfloor M/2\rfloor-1\}$ and the last half $S_{1}=\{q_{j};~{}\lfloor M/2\rfloor\leq j\leq M-1\}$. By applying $P$ to the modes in $S_{1}$ at $t=T/2$, the sign of $H_{j,k}$ for $\{j,k;~{}q_{j}\in S_{0}$,  $q_{k}\in S_{1}\}$ flips and the phonon hopping between the two sets occurring in the later half period cancels with that occurred in the first half period, i.e., $q_{j}\in S_{0}$ and $q_{k}\in S_{1}$ are decoupled. Then, in the second step, we split $S_{0}$ into $S_{00}$ and $S_{10}$ in the same manner as above and apply $P$ to the modes in $S_{10}$ at $t=T/4$ and $43T/$ 4, which flips the sign of $H_{j,k}$ for $\{j,k;~{}q_{j}\in S_{00}$,  $q_{k}\in S_{10}\}$. Similarly, we split $S_{1}$ into $S_{01}$ and $S_{11}$, and apply $P$ to the modes in $S_{11}$ at $t=T/4$ and $43T/$ 4.
[…]
Finally, we need to compensate the unwanted $\pi$-phase shift. For the two-mode DD, $P_{j}$ in $U_{\mathrm{DD}}$ can be moved to the left using the relation $B^{\mathrm{TM}}_{1,0}(\theta)P_{j}=P_{j}B^{\mathrm{TM}}_{1,0}(-\theta)$ as shown by Eq. (II.1.1). For the $M$-mode DD, we can move all the $P_{j}$ in $U_{\mathrm{DD}}$ to the left in a similar manner so that we have $P_{j}^{\Pi_{j}}$, which is equal to $I$ when $\Pi_{j}$ is even and is equal to $P_{j}$ when $\Pi_{j}$ is odd. Because $\Pi_{j}$ is odd if the $q_{j}$ is in $S_{1}$, we apply $P$ to all the modes in $S_{1}$ at $t=T$. Therefore, after applying this compensation procedure, the number of $P_{j}$ is given by $\Pi_{j}=\alpha_{j}+\delta_{1,r}$, where $r$ is the right-most bit of $\alpha_{j}$ and $\delta_{1,r}$ is the Kronecker delta.
Because $P$ is not applied to $q_{0}$, which corresponds to the binary number $\alpha_{0}=00\dots 0$, we can immediately implement $B_{\mathcal{S}}$ defined by Eq. (5) using the DD described above. By using the fact that $P$ is not applied to $q_{0}$, which is associated with the binary number $\alpha_{0}=00\dots 0$, the DD described above can be immediately used to implement $B_{\mathcal{S}}$ defined by Eq. (5). First, we define $M=|\mathcal{S}^{c}|+1$ and choose a mode $q^{\prime}$ from $\mathcal{S}$. Then, by regarding $q^{\prime}$ as the zeroth mode, we apply the $M$-mode DD to the set of modes $\mathcal{S}^{\prime}=\{q^{\prime}\}\cup\mathcal{S}^{c}$. Because the phonon hoppings among the modes in $\mathcal{S}^{c}$ and that between $q^{\prime}$ and any mode in $\mathcal{S}^{c}$ are cancelled and because the choice of $q^{\prime}$ to which $P$ is not applied is arbitrary, the phonon hopping is not affected only among $\mathcal{S}$ by the DD, and consequently, $B_{\mathcal{S}}$ is securely realized.
We note that the present algorithm also works even if the roles of the subscripts ”0” and ”1” are interchanged at each step even though we assign the subscript ”1” to the subsets to which $P$ is applied to obtain the simple relation between $\alpha_{j}$ and $\Pi_{j}$. For example, at $l$-th step, we can apply $P$ to the modes in $S_{0\alpha_{l-1}}$ instead of the modes in $S_{1\alpha_{l-1}}$. As shown in III.2.1, this degree of freedom can be exploited to reduce the error in the DD.
The unitary operator $U_{\mathrm{DD}}$ describing the DD consists of the repetition of the set of $B_{M}(\bm{\theta}_{\tau})$ and $P$ as Eqs. (8) and (10). The number $N_{\mathrm{BP}}$ of this set in $U_{\mathrm{DD}}$ is given by $N_{\mathrm{BP}}=2^{l_{\max}}$. In practice, considering the finite pulse duration $T_{P}$ of $P$, $N_{\mathrm{BP}}$ is upper bounded by $N_{\mathrm{BP}}<T/T_{P}$. Therefore, when $T$ and $T_{P}$ are given, the upper bound for $M$ is obtained as $\lceil\log_{2}M\rceil<\log_{2}(T/T_{P})\Rightarrow$ $M<2^{\lfloor\log_{2}(T/T_{P})\rfloor}$, which limits the applicability of the DD to a large $M$.
The problem on the applicability of the DD to the large $M$ can be addressed by truncating the Coulomb interaction at a certain mode separation $\eta\in\mathbb{Z}$, i.e., by omitting $H_{j,k}$ for $|j-k|>\eta$. This truncation can be rationalized because the hopping rate scales as $\kappa_{j,k}\propto(d|j-k|)^{-3}$, where $d$ denotes the distance between two neighboring ions.
For example, when $M=3\eta$, we split them into three sets $S_{0}$, $S_{1}$, and $S_{2}$, where each set contains $\eta$ modes. Because $|j-k|>\eta$ is satisfied for $\{j,k;~{}q_{j}\in S_{0},\,q_{k}\in S_{2}\}$, we do not need to consider the DD to decouple $S_{0}$ and $S_{2}$. First, we apply $P$ to the modes in $S_{1}$ at $t=T/2$, which decouples $q_{j}\in S_{0}$ and $q_{k}\in S_{2}$ form $q_{l}\in S_{1}$. Then, we split each of $S_{0}$, $S_{1}$, and $S_{2}$ into two subsets and repeat the procedure, so that $N_{\mathrm{BP}}$ is given by $2^{\beta}$ with $\beta=\mathrm{min}(\lceil\log_{2}\eta\rceil+1,\,l_{\max})$.
II.1.3 Repeated-DD
As shown by Eq. (8), the two-mode DD can completely cancel the phonon hopping, but the $M$-mode DD with $M\geq 3$ cannot do so even if $T_{P}=0$. Because we associate ”1” in the subscript of $S_{\alpha_{j}}$ with the application of $P$, the unitary operator describing the 3-mode DD is given by
$U_{\mathrm{DD}}=P_{2}P_{1}B_{3}(\bm{\theta}_{\tau})P_{2}B_{3}(\bm{\theta}_{\tau})P_{2}P_{1}B_{3}(\bm{\theta}_{\tau})P_{2}B_{3}(\bm{\theta}_{\tau}),$ (10)
where $\tau=T/N_{\mathrm{BP}}$ and $N_{\mathrm{BP}}=4$. By introducing the notation
$B_{\pm\pm\pm}$ $\equiv\exp\left[-\mathrm{i}\frac{\tau}{2}\left(\pm\kappa_{1,0}a_{1}^{\dagger}a_{0}\pm\kappa_{2,1}a_{2}^{\dagger}a_{1}\pm\kappa_{2,0}a_{2}^{\dagger}a_{0}\right.\right.$
$\left.\left.\hskip 170.71652pt+\mathrm{h.c.}\right)\right],$ (11)
and using the relations (see Appendix B)
$B_{\pm\pm\pm}P_{1}=P_{1}B_{\mp\mp\pm},$ (12)
$B_{\pm\pm\pm}P_{2}=P_{2}B_{\pm\mp\mp},$ (13)
and $P_{j}^{2}=I$, Eq. (10) can be rewritten as
$U_{\mathrm{DD}}$ $=P_{2}P_{1}B_{+++}P_{2}B_{+++}P_{2}P_{1}B_{+++}P_{2}B_{+++}$
$=B_{-+-}B_{--+}B_{+--}B_{+++},$ (14)
where $B_{3}=B_{+++}$. Unlike the 2-mode DD (8), the four operators in the last line of Eq. (II.1.3) do not cancel with each other, and consequently, $U_{\mathrm{DD}}=I$ cannot be satisfied.
We give a rough estimation of the error in the DD by applying the first-order product approximation [16] to $B_{\pm\pm\pm}$ as
$B_{\pm\pm\pm}$ $=B^{\mathrm{TM}}_{1,0}(\pm(\bm{\theta}_{\tau})_{1,0})B^{\mathrm{TM}}_{2,1}(\pm(\bm{\theta}_{\tau})_{2,1})B^{\mathrm{TM}}_{2,0}(\pm(\bm{\theta}_{\tau})_{2,0})$
$\hskip 14.22636pt+\mathcal{O}(\tilde{\alpha}_{\mathrm{comm}}T^{2}N_{\mathrm{BP}}^{-2}),$ (15)
where $\tilde{\alpha}_{\mathrm{comm}}=\sum_{\xi,\xi^{\prime}=\{(1,0),(2,1),(2,0)\}}(-\kappa_{\xi}\kappa_{\xi^{\prime}}/4)\|[\alpha_{\xi},\alpha_{\xi^{\prime}}]\|$, $\alpha_{\xi=(j,k)}=a_{j}^{\dagger}a_{k}+a_{j}a_{k}^{\dagger}$, and $\|A\|\geq 0$ stands for the spectral norm of $A$. Due to the relations $[a_{j},a_{k}]=0$ for all $\{j,k\}$ and $[a^{\dagger}_{j},a_{k}]=0$ for $j\neq k$, $[\alpha_{(j,k)},\alpha_{(j^{\prime},k^{\prime})}]$ becomes non-zero only when $(j=j^{\prime},k\neq k^{\prime})$ or $(j\neq j^{\prime},k=k^{\prime})$ is satisfied. Because both of $B^{\mathrm{TM}}_{j,k}(+(\bm{\theta}_{\tau})_{j,k})$ and $B^{\mathrm{TM}}_{j,k}(-(\bm{\theta}_{\tau})_{j,k})$ appear twice and cancel with each other in the last line of Eq. (II.1.3), we can obtain $U_{\mathrm{DD}}=I+\mathcal{O}(\tilde{\alpha}_{\mathrm{comm}}T^{2}N_{\mathrm{BP}}^{-2})$. Furthermore, because the coupling strength $\kappa_{j+1,j}$ becomes the largest for the nearest-neighbor modes and is given by $\kappa_{j+1,j}=\kappa_{1,0}$ for an equidistant chain of ions, $\tilde{\alpha}_{\mathrm{comm}}$ can be approximated by $(-\kappa_{1,0}^{2}/4)\sum_{\xi,\xi^{\prime}}\|[\alpha_{\xi},\alpha_{\xi^{\prime}}]\|$, where $\xi=(j+1,j)$ and $\xi^{\prime}=(j^{\prime}+1,j^{\prime})$, and consequently, the error can be rewritten as $\mathcal{O}(\kappa_{1,0}^{2}T^{2}N_{\mathrm{BP}}^{-2})$. It should be noted here that the time scale $T$ of the phonon hopping between the nearest-neighbor modes is proportional to $\kappa_{1,0}^{-1}$, e.g., the 50:50 beam-splitter between the nearest-neighbor modes can be realized with $T=(\pi/4)/(\kappa_{1,0}/2)$. Therefore, for inducing a fixed amount of phonon hopping, $\kappa_{1,0}^{2}$ cancels $T^{2}$ in the error term $\mathcal{O}(\kappa_{1,0}^{2}T^{2}N_{\mathrm{BP}}^{-2})$. Moreover, because $N_{\mathrm{BP}}$ is determined by the algorithm explained in II.1.2, there is no free parameter which we can use to reduce the error in the DD during a fixed amount of phonon hopping.
In order to reduce the error, we replace $\tau$ in Eq. (10) by $\tau/n_{r}$ with a positive integer $n_{r}$, and repeat the pulse sequence $n_{r}$ times as
$\tilde{U}_{\mathrm{DD}}^{n_{r}}$ =$=\left\{P_{2}P_{1}B_{3}\left(\bm{\theta}_{\tau/n_{r}}\right)P_{2}B_{3}\left(\bm{\theta}_{\tau/n_{r}}\right)\right.$
$\left.\hskip 17.07164ptP_{2}P_{1}B_{3}\left(\bm{\theta}_{\tau/n_{r}}\right)P_{2}B_{3}\left(\bm{\theta}_{\tau/n_{r}}\right)\right\}^{n_{r}},$ (16)
which we call a repeated-DD. When $B_{3}$ is approximated by the product of $B^{\mathrm{TM}}_{j,k}((\bm{\theta}_{\tau})_{j,k}/n_{r})$ as in Eq. (II.1.3), the error, which is scaled as in $\tilde{U}_{\mathrm{DD}}^{n_{r}}=I+\mathcal{O}(\tilde{\alpha}_{\mathrm{comm}}T^{2}N_{\mathrm{BP}}^{-2}n_{r}^{-1})$, can be approximated to be $\mathcal{O}(\kappa_{1,0}^{2}T^{2}N_{\mathrm{BP}}^{-2}n_{r}^{-1})$. Therefore, by increasing $n_{r}$, we can suppress the error and obtain $\tilde{U}_{\mathrm{DD}}^{n_{r}}\simeq I$.
II.2 Non-perturbative modulation of a harmonic potential
An off-resonant sideband pulse can be used to implement the $\pi$-phase shift, which can be applied to any number of phonons [7]. However, its interaction rate has been estimated to be less than $2.5\,\mathrm{kHz}$ in Ref. [15], which means that it would take $200\,\mathrm{\mu s}$ to achieve a $\pi$-phase shift. The duration can be reduced to a few tens of microseconds using a resonant sideband pulse [15], but this scheme is only applicable when the phonon number is 0 or 1.
Lau and James [14] has shown that a phase shift which is faster and applicable to any number of phonons can be constructed by modulating the trap potential instead of irradiating the ions with laser light. We describe here this trap-potential modulation approach for a single harmonic oscillator whose Hamiltonian is given by $h_{0}=\hbar\omega_{0}(a^{\dagger}a+1/2)$. As shown in Appendix C, when we apply a quadratic potential,
$H^{\mathrm{LJ}}(t)=\frac{\hbar\Omega^{2}(t)}{4\omega_{0}}\left(a^{\dagger}+a\right)^{2},$ (17)
the total Hamiltonian $h_{0}+H^{\mathrm{LJ}}(t)$ can be interpreted in the coordinate space as the harmonic oscillator oscillating with the time-dependent frequency $\omega(t)$, where $\Omega^{2}(t)=\omega^{2}(t)-\omega_{0}^{2}$. When $\omega(t)$ is related with a real-valued and continuous function $b(t)$ as
$\omega(t)=\sqrt{\left(\frac{\omega_{0}^{2}}{b^{3}}-\ddot{b}\right)/b},$ (18)
and the boundary conditions $b(t\leq 0)=b(t\geq T_{P})=1$ and $\omega(0)=\omega(T_{P})=\omega_{0}$ are satisfied, the application of $H^{\mathrm{LJ}}(t)$ from $t=0$ to $T_{P}$ introduce a phase-shift operator, $\exp(-\mathrm{i}\phi(a^{\dagger}a+1/2))$, with the amount of the phase shift given by
$\phi=\omega_{0}\left(\int_{0}^{T_{P}}\frac{dt}{b^{2}}-T_{P}\right).$ (19)
In addition, the $b$ function should be chosen so that the argument in the square root of the right-hand side of Eq. (18) is always non-negative and that $\ddot{b}$ is continuous. We denote the $\pi$-phase shift induced by $H^{\mathrm{LJ}}$ as $P^{\mathrm{LJ}}$. By replacing $P$ in $U_{\mathrm{DD}}$ by $P^{\mathrm{LJ}}$, we can cancel the Coulomb coupling by the modulation of harmonic potential. Hereafter, we call this approach the cancellation of Coulomb coupling by modulating harmonic potential (C3PO). We note that $P^{\mathrm{LJ}}=\exp(-\mathrm{i}\pi(a^{\dagger}a+1/2))$ differs from the conventional definition of the $\pi$-phase shift (7) by the factor $\exp(-\mathrm{i}\pi/2)$, which only affects the global phase. We adopt the $b$ function defined by using the error function as
$b=\left\{\begin{array}[]{cc}1-\frac{k}{2}\left(1+\mathrm{erf}\left[\left(\frac{t}{T_{\mathrm{u}}}-\frac{1}{2}\right)\sigma\right]\right)&(0\leq t\leq T_{\mathrm{u}})\\ 1-k&(T_{\mathrm{u}}<t<T_{P}-T_{\mathrm{d}})\\ 1-\frac{k}{2}\left(1-\mathrm{erf}\left[\left(\frac{t-(T-T_{\mathrm{d}})}{T_{\mathrm{d}}}-\frac{1}{2}\right)\sigma\right]\right)&(T_{P}-T_{\mathrm{d}}\leq t\leq T_{P})\end{array}\right.,$ (20)
where $T_{\mathrm{u}}$ and $T_{\mathrm{d}}$ are the ramp-up and the ramp-down time of the potential modulation, respectively. The width of the error function $\sigma$ should be large enough to ensure that $b$ is smooth at $t=T_{\mathrm{u}}$ and $T-T_{\mathrm{d}}$. In the present study, we fix $\sigma=6$ and $T_{\mathrm{u}}=T_{\mathrm{d}}=T_{\mathrm{ud}}$. For a given pulse duration $T_{P}$, the strength $k$ is determined so that Eq. (19) is satisfied for $\phi=\pi$ (see Appendix C).
Because the secular frequency of an ion trapped in a radio-frequency (RF) trap is dependent on the amplitude of the RF and the DC voltages [17], the time-dependent frequency $\omega(t)$ can be implemented by the modulation of the RF and/or DC voltage(s). The relation between $\omega(t)$ and the RF/DC amplitudes for a linear trap is given in Appendix D. However, it is not an easy task to control the trap potential for one ion while keeping the potentials for the other ions the same. Using a surface electrode trap, it was shown that the secular frequency of a local mode can be adjusted without affecting the other modes by applying a specific set of DC voltages [18]. Therefore, it is expected that the C3PO, which requires the application of $P^{\mathrm{LJ}}$ by individually modulating the trap frequency, can be implemented using a surface electrode trap.
III Results
We demonstrate the C3PO by numerically solving the time-dependent Schrödinger equation (TDSE) from $t=0$ to $T$, where $T$ is set to the period of the 50:50 beam-splitter between the neighbouring modes. In order to solve the TDSE efficiently and accurately, the semi-global method [19] has been adopted. There are two sources of error in the C3PO; (i) the algorithmic error due to the imperfect cancellation of the phonon hopping as explained in II.1.2 and (ii) the gate error due to the finite pulse duration of $P^{\mathrm{LJ}}$. Because the algorithmic error is absent for $M=2$, we first examine the 2-mode C3PO to determine the optimal parameters which minimize the gate error. Then, both the algorithmic and the gate error are investigated for $M=3$. Hereafter, when the instantaneous $\pi$-phase shift is applied instead of $P^{\mathrm{LJ}}$, we call the DD the ideal DD. We evaluate the error by $\braket{E}\equiv|\bra{\psi_{0}}I-U\ket{\psi_{0}}|=1-|\bra{\psi_{0}}U\ket{\psi_{0}}|$ | for an initial state |$\ket{\psi_{0}}$, where $U$ represents the pulse sequence of the ideal DD or the C3PO.
III.1 2-mode C3PO
We consider a chain of $\mathrm{{}^{40}Ca^{+}}$ ions. The oscillation frequency along $\tilde{x}$ is set to $\omega_{0}/(2\pi)=2.2\,\mathrm{MHz}$. The state of $M$ phonons is represented as |$\ket{\psi}=\sum_{j}c_{j}\ket{j}$, where |$\ket{j}=\ket{n_{M-1},\dots,\,n_{1},\,n_{0}}$ and $n_{m}$ is the number of phonons in the $m$-th mode.
[…]
Figure 1: The populations of a 2-mode phonon state during the C3PO with $d=27.6\,\mathrm{\mu m}$. (a) $T_{P}=4.0\,\mathrm{\mu s}$ and (b) $T_{P}=1.0\,\mathrm{\mu s}$. The initially occupied phonon state is shown by the solid curve, while the other states are shown by the dashed curves. The application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas.
Figure 2: The populations of a 2-mode phonon state during the C3PO. $T_{P}=4.0\,\mathrm{\mu s}$ and $d=43.8\,\mathrm{\mu m}$. The initially occupied phonon state is shown by a solid curve while the other states are shown by dashed curves. The application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas.
We can reduce the error by shortening the pulse duration or/and slowing down the phonon hopping. For example, as shown in Fig. 1 (b), when $T_{P}$ is set to $1\,\mathrm{\mu s}$, the error is lowered to be $\braket{E}=2.2\times 10^{-6}$. Instead of making $T_{P}$ four times smaller, which requires an improvement in the experimental setup, we may make the phonon hopping slower by four times, which can be easily achieved by replacing $d$ by $4^{(1/3)}d$. As shown in Fig. 2, for $d=43.8\,\mathrm{\mu m}$, which corresponds to $\kappa_{1,0}/(2\pi)=0.47\,\mathrm{kHz}$, the error is lowered further to $\braket{E}=2.2\times 10^{-8}$.
Although we only show the populations of the states that preserve the number of phonons, other states can also be populated during the application of $P^{\mathrm{LJ}}$. This can be most clearly seen in the case of $T_{P}=1\,\mathrm{\mu s}$. The expanded view of the shaded area of Fig. 1 (b) is given in Fig. 7 in Appendix C.
III.2 3-mode C3PO
[…]
III.2.1 Ideal DD
Figure 3: The populations of a 3-mode phonon state during the single cycle of the ideal DD. The pulse sequecne is defined by Eq. (10). The initially occupied phonon state is shown by a solid curve while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines.
[…]
$U_{\mathrm{DD}}$ $=P_{2}P_{1}B_{3}(\bm{\theta}_{\tau})P_{1}B_{3}(\bm{\theta}_{\tau})P_{2}P_{1}B_{3}(\bm{\theta}_{\tau})P_{1}B_{3}(\bm{\theta}_{\tau})$ (21)
$=B_{-+-}B_{+--}B_{--+}B_{+++},$ (22)
and the error is reduced to $\braket{E}=4.4\times 10^{-5}$. This significant improvement shows the importance of the optimization of the pulse sequence, which will be discussed further in a future work.
Although the pulse sequence is designed so that the initial state is recovered at $t=T$, the initial state is almost recovered at $t=T/2$ in Fig. 4 (a), and the dynamics from $t=0$ to $T/2$ is repeated from $t=T/2$ to $T$. This is because $P_{1}$ applied at $t=T/4$ flips the sign of the nearest neighbour coupling terms, $a_{1}^{\dagger}a_{0}+\mathrm{h.c.}$ and $a_{2}^{\dagger}a_{1}+\mathrm{h.c.}$, which have the dominant contribution to the dynamics, while $P_{2}$ applied at $t=T/2$ flips the sign of the non-nearest neighbour coupling term $a_{2}^{\dagger}a_{0}+\mathrm{h.c.}$, whose contribution is eight times smaller in magnitude because $\kappa_{2,0}\propto(2d)^{-3}$. This observation supports the validity of the truncation described in II.1.2, where the pulse sequence is constructed to cancel the hopping between ions in the distance smaller than or equal to $\eta d$.
The ideal repeated-DD $\tilde{U}_{\mathrm{DD}}^{n_{r}}$ derived from Eq. (21) can be defined similarly as in Eq. (II.1.3) and an example with $n_{r}=5$ is illustrated in Fig. 5 (a). We can see the five cycles of phonon hopping and its cancellation. We note that the error scaling given in II.1.3 is not tight and the actual scaling for fixed $\kappa_{1,0}$ and $N_{\mathrm{BP}}$ is estimated as $\mathcal{O}(T^{4}n_{r}^{-2})$ from the numerical simulation of the 3-mode DD. Here, the error for the repeated-DD with $n_{r}=5$ is calculated to be $\braket{E}=1.9\times 10^{-6}$, which is $n_{r}^{2}=25$ times smaller than the error obtained with $n_{r}=1$.
Figure 4: The populations of a 3-mode phonon state during the single cycle of (a) the ideal DD and (b) the C3PO. The initially occupied phonon state is shown by a solid curve while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines in (a) while the application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas in (b).
Figure 5: The populations of a 3-mode phonon state during the five cycles of (a) the ideal DD and (b) the C3PO. The initially occupied phonon state is shown by a solid curve while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines in (a) while the application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas in (b).
III.2.2 C3PO
By replacing $P_{j}$ in Eq. (21) by $P^{\mathrm{LJ}}_{j}$, a numerical simulation of the 3-mode C3PO was performed with $n_{r}=1$ and $n_{r}=5$. The results of the simulation with $n_{r}=1$ and $n_{r}=5$ are shown in Fig. 4 (b) and Fig. 5 (b), respectively. The errors obtained for $n_{r}=1$ and $n_{r}=5$ are $\braket{E}=4.4\times 10^{-5}$ and $\braket{E}=2.6\times 10^{-6}$, respectively.
Although the algorithmic error can be reduced further by increasing $n_{r}$ in the case of the ideal DD, the upper limit of $n_{r}$ given using $N_{\mathrm{BP}}n_{r}<T/T_{P}$ exists for the C3PO due to the finite pulse duration. In the current case with $N_{\mathrm{BP}}=4$ and $T/T_{P}\simeq 132$, we obtain $n_{r}<33$. By using $N_{\mathrm{BP}}=2^{\lceil\log_{2}M\rceil}$, the upper bound of $M$ for a fixed $n_{r}$ is given as $M<2^{\lfloor\log_{2}(T/(T_{P}n_{r}))\rfloor}$. For example, $M<128$ for $n_{r}=1$ and $M<16$ for $n_{r}=5$. Therefore, in order to apply the C3PO to a system with a large $M$, we need to adopt the truncation of $H_{j,k}$ for $|j-k|>\eta$. By requiring $N_{M}=2^{\lceil\log_{2}\eta\rceil+1}<T/(T_{P}n_{r})$, we obtain $\eta<2^{\lfloor\log_{2}(T/(T_{P}n_{r}))\rfloor-1}$, which means that, for example, $\eta$ needs to be $\eta<64$ for $n_{r}=1$ and $\eta<8$ for $n_{r}=5$.
IV Conclusion
We have proposed a method called C3PO to control phonon hopping among the local modes in a chain of trapped ions by applying a sequence of phase-shift operations realized by the modulation of the trap potential. The present method can be applied to a system having any number of phonons and modes, which has been demonstrated by the numerical simulations of the cancellation of the phonon hopping among local modes. By applying the C3PO to cancel the phonon hopping among all the modes but the chosen set of modes $\mathcal{S}$, we can implement the multi-mode entangling gate for $\mathcal{S}$. It should be noted that the C3PO can be implemented to the boson sampling [7] and the logical two-qubit gate for the GKP state [20]. For the next step for the experimental demonstration of the C3PO, we need to examine possible sources of error such as imperfect control of the phase shift pulse, motional heating, and the vibrational decoherence. We note that the logical two-qubit gate for the GKP states encoded using two radial modes of a single ion was recently demonstrated [21], but this approach can be applied to a system whose mode number is less than or equal to 3 because it relies on the coupling between the electronic and vibrational degrees of freedom of a single ion.
Acknowledgements.
This work was supported by JST-CREST Grant No. JPMJCR23I7.
Appendix A Hamiltonian
We consider $M$ trapped ions oscillating along $\tilde{x}$ axis. The Hamiltonian of $j$-th ion is given by $H_{j}=\tilde{p}_{j}^{2}/(2m)+m\tilde{\omega}_{0}^{2}\tilde{x}_{j}^{2}/2$, where $m$ is the ion’s mass and $\tilde{\omega}_{0}$ is the oscillation angular frequency. The Coulomb interaction between two cations located at positions $\bm{r}_{1}$ and $\bm{r}_{2}$ can be approximated as [20]
$\frac{e^{2}}{4\pi\epsilon_{0}}\frac{1}{|\bm{r}_{1}-\bm{r}_{2}|}\simeq\frac{e^{2}}{4\pi\epsilon_{0}}\left(\frac{1}{d_{1,2}}+\frac{-\tilde{x}_{1}^{2}-\tilde{x}_{2}^{2}+2\tilde{x}_{1}\tilde{x}_{2}}{2d_{1,2}^{3}}\right),$ (23)
where $d_{1,2}$ is the distance between the two equilibrium positions of the two cations. In deriving Eq. (A1), we use the fact that the displacement of each ion from the equilibrium position is perpendicular to the trap axis. Hereafter, we neglect the first term in the right-hand side of Eq. (A1) because it only affects the global phase. Then, the total Hamiltonian for $M$ harmonic oscillators interacting through the Coulomb interaction is given by
$H$ $=\sum_{j}\left(\frac{\tilde{p}_{j}^{2}}{2m}+\frac{m\tilde{\omega}_{0}^{2}\tilde{x}_{j}^{2}}{2}\right)$
$\hskip 14.22636pt+\sum_{j>k}\left(\frac{-m\tilde{\omega}_{0}\tilde{\kappa}_{j,k}}{2}(\tilde{x}_{j}^{2}+\tilde{x}_{k}^{2})+\frac{\hbar\tilde{\kappa}_{j,k}}{x_{0}^{2}}\tilde{x}_{j}\tilde{x}_{k}\right),$ (24)
where $\tilde{\kappa}_{j,k}=e^{2}/(4\pi\epsilon_{0}d_{j,k}^{3}m\tilde{\omega}_{0})$ and $x_{0}=\sqrt{\hbar/(m\omega_{0})}$. As can be seen in Eq. (A), the first term of the Coulomb interaction shifts the oscillation frequency while the second term mixes the different modes. Because the frequency shift depends on $j$ as represented by $\sum_{k}\tilde{\kappa}_{j,k}$, $P^{\mathrm{LJ}}$ should also depends on $j$, which induces additional complexity.
Therefore, we replace $\tilde{\omega}_{0}$ by the $j$-dependent frequency $\tilde{\omega}_{j}$ in a manner that the shifted oscillation frequency becomes $j$-independent. First, the Hamiltonian can be written as
$H$ $=\sum_{j}\left(\frac{\tilde{p}_{j}^{2}}{2m}+\frac{m\tilde{\omega}_{j}^{2}\tilde{x}_{j}^{2}}{2}+\frac{-m\sum_{k\neq j}\tilde{\omega}_{j,k}\tilde{\kappa}_{j,k}}{2}\tilde{x}_{j}^{2}\right)$
$\hskip 14.22636pt+\sum_{j>k}\frac{\hbar\tilde{\kappa}_{j,k}}{x_{0,j}x_{0,k}}\tilde{x}_{j}\tilde{x}_{k},$ (25)
$x_{0,j}=\sqrt{\hbar/(m\tilde{\omega}_{j})}$, and $\tilde{\kappa}_{j,k}$ is redefined as $\tilde{\kappa}_{j,k}=e^{2}/(4\pi\epsilon_{0}d_{j,k}^{3}m\tilde{\omega}_{j,k})$ with $\tilde{\omega}_{j,k}=\sqrt{\tilde{\omega}_{j}\tilde{\omega}_{k}}$. We eliminate the $j$-dependence of the oscillation frequency by requiring
$\omega_{0}^{2}$ $=\tilde{\omega}_{j}^{2}-\sum_{k\neq j}\tilde{\omega}_{j,k}\tilde{\kappa}_{j,k}$
$=\tilde{\omega}_{j}^{2}-\sum_{k\neq j}\frac{e^{2}}{4\pi\epsilon_{0}d_{j,k}^{3}m}.$ (26)
Therefore, for a given $\omega_{0}$, by tuning the harmonic potential of each mode so that the bare oscillation frequency $\tilde{\omega}_{j}$ satisfies Eq. (26), Eq. (A) can be rewritten as
$H$ $=\sum_{j}\left(\frac{\tilde{p}_{j}^{2}}{2m}+\frac{m\omega_{0}^{2}\tilde{x}_{j}^{2}}{2}\right)+\sum_{j>k}\frac{\hbar\tilde{\kappa}_{j,k}}{x_{0,j}x_{0,k}}\tilde{x}_{j}\tilde{x}_{k}.$ (27)
Then, by introducing the scaled coordinate and momentum defined respectively as $x=\tilde{x}/x_{0}$ and $p=\tilde{p}/p_{0}$ with $x_{0}=\sqrt{\hbar/(m\omega_{0})}$ and $p_{0}=\sqrt{\hbar m\omega_{0}}$, the Hamiltonian is given by
$H=\sum_{j}\frac{\hbar\omega_{0}}{2}(p_{j}^{2}+x_{j}^{2})+\sum_{j>k}\hbar\kappa_{j,k}x_{j}x_{k}.$ (28)
By using the relations, $x_{j}=(a^{\dagger}_{j}+a_{j})/\sqrt{2}$ and $p_{j}=\mathrm{i}(a^{\dagger}_{j}-a_{j})/\sqrt{2}$, the Hamiltonian in the Fock basis is obtained as
$H=$ $\sum_{j}\hbar\omega_{0}\left(a_{j}^{\dagger}a_{j}+\frac{1}{2}\right)$
$+\sum_{j>k}\frac{\hbar\kappa_{j,k}}{2}\left(a_{j}^{\dagger}+a_{j}\right)\left(a_{k}^{\dagger}+a_{k}\right).$ (29)
When the rotating-wave approximation is applied, the terms which do not preserve the number of phonons vanish and the resultant Hamiltonian becomes
$H$ $=H_{0}+\sum_{j>k}H_{j,k}$
$=\sum_{j}\hbar\omega_{0}\left(a_{j}^{\dagger}a_{j}+\frac{1}{2}\right)+\sum_{j>k}\frac{\hbar\kappa_{j,k}}{2}\left(a_{j}^{\dagger}a_{k}+a_{j}a_{k}^{\dagger}\right).$ (30)
Appendix B Phase shift and beam-splitter
The idea of the DD is based on the following relations;
$e^{\mathrm{i}\phi a^{\dagger}a}ae^{-\mathrm{i}\phi a^{\dagger}a}=e^{-\mathrm{i}\phi}a,$ (31)
$e^{\mathrm{i}\phi a^{\dagger}a}a^{\dagger}e^{-\mathrm{i}\phi a^{\dagger}a}=e^{\mathrm{i}\phi}a^{\dagger}.$ (32)
When the two-mode beam-splitter is defined as $B^{\mathrm{TM}}_{1,0}(\theta)=\exp(-\mathrm{i}\theta(a_{1}^{\dagger}a_{0}+a_{1}a_{0}^{\dagger}))$ and the phase shift is defined as $P_{1}(\phi)=\exp(-\mathrm{i}\phi a_{1}^{\dagger}a_{1})$, we derive
$P_{1}^{\dagger}(\phi)B^{\mathrm{TM}}_{1,0}(\theta)P_{1}(\phi)=\exp\left(-\mathrm{i}\theta\left(e^{\mathrm{i}\phi}a_{1}^{\dagger}a_{0}+e^{-\mathrm{i}\phi}a_{1}a_{0}^{\dagger}\right)\right),$ (33)
and the similar equations for $P_{0}(\phi)$ and $P_{0}^{\dagger}(\phi)$. Therefore, by denoting the $\pi$-phase shift as $P_{j}$ and using $P_{j}=P_{j}^{\dagger}$, we derive
$P_{j}B^{\mathrm{TM}}_{1,0}(\theta)P_{j}=B^{\mathrm{TM}}_{1,0}(-\theta),\,j=0~{}\mathrm{or}~{}1.$ (34)
Because the phonon-hopping Hamiltonian satisfies
$P_{\xi}H_{j,k}P_{\xi}=-H_{j,k},\,\xi=j~{}\mathrm{or}~{}k,$ (35)
Eq. (34) can be extended to the $M$-mode entangling gate $B_{M}(\bm{\theta})=\exp(-\mathrm{i}\tau\sum_{j>k}H_{j,k}/\hbar)=\exp(-\mathrm{i}\sum_{j>k}\theta_{j,k}(a_{j}^{\dagger}a_{k}+a_{j}a_{k}^{\dagger}))$ as
$P_{j}B_{M}(\bm{\theta})P_{j}$
$\hskip 28.45274pt=\exp\left(-\mathrm{i}\sum_{j^{\prime}>k}(-1)^{\delta_{j,j^{\prime}}}\theta_{j^{\prime},k}\left(a_{j^{\prime}}^{\dagger}a_{k}+a_{j^{\prime}}a_{k}^{\dagger}\right)\right),$ (36)
where $\delta_{j,j^{\prime}}$ is the Kronecker delta. Consequently, because $P_{j}$ flips the sign of the phonon hopping terms containing $a_{j}$ or $a_{j}^{\dagger}$, Eqs. (12) and (13) are derived.
Appendix C Time-dependent harmonic oscillator
A time-dependent harmonic oscillator (TDHO) is defined by the following Hamiltonian
$H=\frac{\tilde{p}^{2}}{2m}+\frac{1}{2}m\omega^{2}(t)\tilde{x}^{2},$ (37)
which can be decomposed into $h_{0}=\tilde{p}^{2}/(2m)+m\omega_{0}^{2}\tilde{x}^{2}/2$ and the time-dependent modulation
$H^{\mathrm{LJ}}(t)$ $=H-h_{0}=\frac{1}{2}m\Omega^{2}(t)\tilde{x}^{2}$
$=\frac{\hbar\Omega^{2}(t)}{2\omega_{0}}x^{2}$
$=\frac{\hbar\Omega^{2}(t)}{4\omega_{0}}(a^{\dagger}+a)^{2},$ (38)
where $\Omega^{2}(t)=\omega^{2}(t)-\omega_{0}^{2}$. The solution of $\mathrm{i}\partial_{t}\ket{\psi}=H\ket{\psi}$ is known to be given by a closed form [22, 23]. When the initial state at $t=0$ is the $n$-th eigenstate $\varphi_{n}$ of $h_{0}$, the solution at $t$ is given in the scaled coordinate by
$\psi=\sqrt{\frac{1}{2^{n}n!\sqrt{\pi}bx_{0}}}\mathcal{H}_{n}\left(\frac{x}{b}\right)\exp\left(-\frac{1}{2}\left(\frac{x}{b}\right)^{2}\right)\exp\left(\frac{\mathrm{i}\dot{b}}{2b\omega_{0}}x^{2}\right)\exp\left(-\mathrm{i}\left(n+\frac{1}{2}\right)\omega_{0}\int_{0}^{T_{P}}\frac{dt^{\prime}}{b^{2}}\right),$ (39)
where $\mathcal{H}_{n}$ is the Hermite polynomial and $b$ is related with $\omega(t)$ by
$\ddot{b}+\omega^{2}(t)b-\frac{\omega_{0}^{2}}{b^{3}}=0.$ (40)
When the boundary conditions, $b(t\leq 0)=b(t\geq T_{P})=1$ and $\omega(0)=\omega(T_{P})=\omega_{0}$, are satisfied, $\psi(t\geq T_{P})$ coincides with $\varphi_{n}$ except the $n$-dependent phase factor. In the interaction picture with respect to $h_{0}$, $\psi$ acquires the phase $\exp(-\mathrm{i}(n+1/2)\phi)$ at $t\geq T_{P}$, where $\phi$ is defined by Eq. (19).
[…]
During the application of $P^{\mathrm{LJ}}$, the phonon number is not necessarily to be preserved. For example, during the application of $P^{\mathrm{LJ}}$ with $T_{P}=1.0\,\mathrm{\mu s}$ in Fig. 1 (b), the populations of the phonon-number-preserving states decrease while the states which do not preserve the phonon number are populated as shown in Fig. 7.
Figure 6: The time-dependent frequency $\omega(t)$ and the $b$ function for the long pulse ($T_{P}=4.0\,\mathrm{\mu s}$) with $T_{\mathrm{ud}}=4.4T_{0}=2.0\,\mathrm{\mu s}$ and $k=0.0529$.
[…]
Appendix D Electric pulse to realize the time-dependent harmonic oscillator
When we apply an oscillating voltage $V(t)=U_{0}+V_{0}\cos(\Omega_{r}t)$ between diagonal electrodes of a linear trap, the quadrupole potential is created near the trap axis $\tilde{z}$ [17], where $U_{0}$ and $V_{0}$ are the amplitudes of the DC and RF voltages, respectively, and $\Omega_{r}$ is the RF frequency. The motion of the ion along the radial directions, $\tilde{x}$ and $\tilde{y}$, can be separated into the secular and micro motions when $|a_{\alpha}|,\,|q_{\alpha}|\ll 1\,(\alpha=\tilde{x},\,\tilde{y})$ are satisfied for $a_{\tilde{x}}=-a_{\tilde{y}}=4eU_{0}/(m\Omega_{r}^{2}r_{0}^{2})$ and $q_{\tilde{x}}=-q_{\tilde{y}}=2eV_{0}/(m\Omega_{r}^{2}r_{0}^{2})$, where $r_{0}$ is the distance from the trap axis to the electrodes. The effective potential for the secular motion is given by
$\Phi_{\mathrm{eff}}=\frac{m}{2e}\left(\omega_{0\tilde{x}}^{2}\tilde{x}^{2}+\omega_{0\tilde{y}}^{2}\tilde{y}^{2}\right),$ (41)
where $\omega_{0\alpha}=\sqrt{a_{\alpha}+q_{\alpha}^{2}/2}\,\Omega_{r}/2$.
In addition, we need to apply a DC voltage to the end electrodes to provide confinement along the $\tilde{z}$ direction. Therefore, the 3-dimensional harmonic potential near the trap axis is given as
$\Phi_{\mathrm{eff}}+\Phi_{\tilde{z}}=\frac{m}{2e}\left(\omega_{\tilde{x}}^{2}\tilde{x}^{2}+\omega_{\tilde{y}}^{2}\tilde{y}^{2}+\omega_{\tilde{z}}^{2}\tilde{z}^{2}\right),$ (42)
where $\omega_{\alpha}=\sqrt{\omega_{0\alpha}^{2}-\omega_{\tilde{z}}^{2}/2}$ for $\alpha=\tilde{x},\tilde{y}$.
In order to realize the time-dependent frequency $\omega(t)$, we modulate $U_{0}$ or $V_{0}$. When we modulate the DC amplitude while fixing the RF amplitude, $\omega_{\alpha}$ at $t$ is given as
$\omega_{\alpha}(t)^{2}=\frac{eU_{0}(t)}{mr_{0}^{2}}+\frac{q_{\alpha}^{2}\Omega_{r}^{2}}{8}-\frac{\omega_{\tilde{z}}^{2}}{2}.$ (43)
By denoting $\omega_{\alpha}(t)$ as $\omega(t)$, we obtain the time-dependent DC amplitude for a given $\omega(t)$ as
$U_{0}(t)=\frac{mr_{0}^{2}}{e}\left(\omega^{2}(t)+\frac{\omega_{\tilde{z}}^{2}}{2}-\frac{q_{\alpha}^{2}\Omega_{r}^{2}}{8}\right).$ (44)
Similarly, when we fix the DC amplitude, the time dependent RF amplitude is given by
$V_{0}(t)=\frac{\sqrt{2}m\Omega_{r}r_{0}^{2}}{e}\sqrt{\omega^{2}(t)+\frac{\omega_{\tilde{z}}^{2}}{2}-\frac{a_{\alpha}\Omega_{r}^{2}}{4}}.$ (45)